OpenStudy (anonymous):

find the product: (x^2+7) (x^3-5x-4)

OpenStudy (lgbasallote):

just distribute \[(x^2 \times x^3) + (x^2 \times -5x) + (x^2 \times -4)\]\[+ (7 \times x^3) + (7 \times -5x) + (7 \times -4)\]

OpenStudy (lgbasallote):

just operate on each grouping one by one

OpenStudy (anonymous):

x^5 +2x^3-4x^2-28

OpenStudy (anonymous):

Thank ya'll!!